solving system if equations by substitution
1/2x+2y=12
x-2y=6

Answer by rfer(16322)

x=2y+6
---------------
1/2x+2y=12
1/2(2y+6)+2y=12
y+3+2y=12
3y=9
y=3
---------------
.5x+2*3=12
.5x=6
x=12
----------------